# Satellite Federated Fine-Tuning for Foundation Models in Space Computing Power Networks

### v2

@@ Abstract
Advancements in artificial intelligence (AI) and low-earth orbit (LEO) satellites have promoted the application of large remote sensing foundation models for various downstream tasks. However, direct downloading of these models for fine-tuning on the ground is impeded by privacy concerns and limited bandwidth. Satellite federated learning (FL) offers a solution by enabling model fine-tuning directly on-board satellites and aggregating model updates without data downloading. Nevertheless, for large foundation models, the computational capacity of satellites is insufficient to support effective on-board fine-tuning in traditional satellite FL frameworks. To address these challenges, we propose a satellite-ground collaborative federated fine-tuning framework. The key of the framework lies in how to reasonably decompose and allocate model components to alleviate insufficient on-board computation capabilities. During fine-tuning, satellites exchange intermediate results with ground stations or other satellites for forward propagation and back propagation, which brings communication challenges due to the special communication topology of space transmission networks, such as intermittent satellite-ground communication, short duration of satellite-ground communication windows, and unstable inter-orbit inter-satellite links (ISLs). To reduce transmission delays, we further introduce tailored communication strategies that integrate both communication and computing resources. Specifically, we propose a parallel intra-orbit communication strategy, a topology-aware satellite-ground communication strategy, and a latency-minimalization inter-orbit communication strategy to reduce space communication costs. Simulation results demonstrate significant reductions in training time with improvements of approximately 33%.
@@ I Introduction
A promising solution to this challenge is satellite federated learning (SFL), which enables satellites to collaboratively train AI models without requiring raw data to be transmitted to ground stations [10, 11, 12]. While SFL has been explored in prior studies, most existing work focuses on small-scale models. Despite the rapid development of Space-CPN, onboard computation capabilities remain significantly constrained compared to terrestrial AI clusters, making the fine-tuning of large models a major challenge. Satellite federated learning, as a solution for distributed training, supports collaborative training among satellites [10, 11]. To solve this problem, it is a feasible solution to decompose and deploy the computing tasks of fine-tuning large models in Space-CPN and use the distributed computation capabilities for federated fine-tuning [13, 14]. According to the distribution of computation capabilities in Space-CPN, the modules are deployed as follows: the embedding layer is deployed on satellites because satellites are data sources, the backbone network is deployed on the terrestrial server, and the classification layer generating outputs is deployed on satellites, as illustrated in Fig. 1.
In this deployment framework, the primary challenge lies in the communication process. During forward propagation, satellites compute local embedding vectors and perform intra-orbit aggregation. Notably, while raw satellite data can reach gigabyte levels in total size, the transmitted embedding vectors for each sample are only a few megabytes, significantly alleviating communication overhead. Once computed, these embedding vectors are transmitted to the ground for feature vector computation. In this process, satellite-ground communication faces some special challenges compared with transmission on the ground [14]. Firstly, due to the movement of satellites, especially LEO satellites that circle the Earth at high speeds, SGLs are dynamic and unstable. Secondly, it takes much time to access the GS for each satellite. Besides, the insufficient bandwidth of SGLs leads to slower data rates and higher latency. Then, feature vectors are sent back to satellites for output generation, which faces similar questions as embedding vector downloading. Finally, satellites update local models and aggregate a global model through ISLs. The dynamic ISLs make link selection and allocation extremely complex in terms of network management.
@@ II-A Centralized Satellite FL
In centralized satellite FL, a coordinator, such as a PS, collects model updates from satellites. Satellites perform local training on their datasets and then send the model updates to the central coordinator. Some algorithms employ only SGLs during training, where models are trained on-board and transmitted via SGLs for global model aggregation at the PS. These algorithms are usually designed to address the challenges of time-varying and unstable SGLs. For instance, the algorithm proposed in [15] adjusted aggregation intervals based on the density of satellite-ground connections to increase the frequency of model updates. [16] proposed an over-the-air computation based scheme and optimized both uplink and downlink communications through beamforming. Considering the variations in connection time and connection density among satellites, asynchronous FL is introduced to mitigate the effects of stragglers, allowing satellites to send model updates to the server asynchronously [17, 18]. A compensation mechanism proposed in [19] addressed gradient staleness by adjusting the updates according to discrepancies.
@@ III-B1 Ground Communication Network
During federated fine-tuning, training information is exchanged between satellites and GSs for global aggregation. As depicted in Fig. 2, there are $G$ GSs and a PS situated on the ground, and all GSs are connected with the PS. The PS serves as the central coordinator in this system, while the GSs act as relays that facilitate communication between the satellites and the PS. The communication links between the GSs and the PS are fixed and characterized by high-speed fabric connectivity, ensuring stable and fast transmission [30]. To represent the characteristics of the ground communication network, we model it as a stable star network, as illustrated in Fig. 3(b), where the PS is the central node and transmission between the central node and other nodes is fast.
@@ III-B3 Satellite-Ground Communication Network
When analyzing SGLs, we sample some time points discretely $\mathcal{T}=\left\{T_{0},T_{1},T_{2},...\right\}$ over a period. A satellite and a GS can communicate between $T_{i}$ and $T_{i+1}$ if at least one link is feasible at $T_{i}$. The interval between two adjacent time points is 1 minute. At different time points, the topology of SGLs varies. Thus, we model satellite-ground communication links as time-varying topology, as depicted in Fig. 3(c).
@@ III-C Satellite Federated Fine-Tuning for Remote Sensing Foundation Models
Our primary objective is to address the computing power constraints of fine-tuning FMs on satellites. To achieve this, we employ a computation task decomposition technique, effectively splitting the FM into distinct components and deploying them based on requirements. To protect data privacy, we deploy the embedding layer and the task-specific head on the satellites. The backbone network, which bears most of the computational workload (approximately 99%), is deployed on the ground. This significantly reduces the computational burden on the satellites, allowing them to handle only about $1\text{\textperthousand}$ of the overall computational tasks. During the forward propagation process, embedding vectors and feature vectors are exchanged between the satellites and the ground. Notably, the size of these vectors is approximately 20 MB for a sample, a reduction of about a thousand times compared to the raw satellite imagery. This drastic reduction in data size not only enhances communication efficiency but also alleviates bandwidth requirements, making the system more cost-effective and practical for deployment in space-constrained environments.
@@ IV-A1 Intra-Orbit Embedding Vector Broadcast
First, each satellite $S_{p,n}$ computes embedding vectors $\bm{E}_{p,n}$ utilizing data stored locally. Then, all computed embedding vectors are broadcast within each orbit so that any satellite connected to a GS can transmit intra-orbit data to the ground for forward propagation. We visualize the connection windows between satellites and GSs throughout the day in Fig. 5 for comparison. Each connection window lasts for approximately $5\sim 10$ minutes. If data transmission starts at 23:30, some satellites establish their first satellite-ground communication around three hours later. So the conventional method, where each satellite only sends its local embedding vectors without coordination, wastes the communication resources of other satellites. In contrast, our proposed method allows other intra-orbit satellites to transmit data immediately without waiting.
Due to the instability of inter-orbit ISLs and the volume of intra-orbit embedding vectors, embedding vectors are not exchanged between different orbits. To optimize this process, we have designed a parallel intra-orbit broadcast strategy, detailed in Section V-A, to make full use of each satellite’s communication capabilities. After the broadcast, each satellite stores the concatenated intra-orbit embedding vectors, represented as:
@@ IV-A2 Embedding Vectors Transmission
Intra-orbit embedding vectors are transmitted from satellites to the PS in each orbit. A key challenge is that satellite-ground communication topology is time-varying, as illustrated in section III-B. To address this, we design a topology-aware satellite-ground transmission strategy in section V-B, which allows multiple orbits to independently utilize this strategy for parallel transmission. The algorithm takes into account the heterogeneity of different SGLs and allocates different transmission tasks to SGLs. Thus, the total transmission time $T_{\text{em}}$ is determined by the maximum transmission time across all orbits:
@@ IV-B3 Inter-orbit Head Aggregation
Head parameters are transmitted from orbit 1 to orbit $P$ through inter-orbit ISLs for global aggregation. To ensure efficient convergence and communication, the inter-orbit ISLs are carefully selected for transmission between two adjacent orbits. A path selection algorithm is proposed in section V-C, which can adapt to dynamic inter-orbit ISLs and automatically select the optimal path to minimize transmission latency.
@@ IV-B4 Global Head Broadcast
Finally, the global model is transmitted back from orbit $P$ to orbit $1$ via inter-orbit ISLs. Within each orbit, the satellite receiving the global head then disseminates it to other satellites for the next round of training, where satellites continue to train using their local data. This step ensures that every satellite has access to the latest version of the global model. Notably, this can be efficiently achieved by simply reversing the direction of transmission during inter-orbit head aggregation.
@@
PS initializes and sends model parameters $\textbf{w}(0)$ to satellites;
for round r=0,…,R  do
@@ V Space Communication Strategies
V-A Intra-Orbit Communication for Partial Head Aggregation
@@
Figure 6: Ring Allreduce based intra-orbit communication. Take N=3 as an example. Satellites $\left\{S_{p,1},S_{p,2},S_{p,3}\right\}$ travel in orbit $p$. In the scatter-reduce phase, data on satellite $S_{p,i}$ is split into $3$ chunks, i.e., $C_{1}^{i},C_{2}^{i}$ and $C_{3}^{i}$. At communication round $1$, $S_{p,1},S_{p,2},S_{p,3}$ transmit $C_{1}^{1},C_{2}^{2},C_{3}^{3}$ respectively. So each satellite only transmits $\frac{1}{3}$ data $2$ times in the scatter-reduce phase and gather phase. The total time is $2\times 2\times\frac{1}{3}\times T$, where $T$ is the time to transmit all data on a satellite.
V-A1 Scatter-Reduce Phase
@@ V-A2 Gather Phase
In the first and second phases of Ring Allreduce, each satellite receives data $N-1$ times, with each transmission involving $\frac{\operatorname{size}(data)}{N}$ bits of data, where $\operatorname{size}(data)$ denotes the total number of bits for size. Thus, the total time is given by:
$T_{\text{intra}}=2\cdot(N-1)\cdot\frac{\operatorname{size}(\text{head})}{N}<2\cdot\operatorname{size}(\text{head}).$ (21)
Ring AllReduce has a constant communication complexity, meaning the communication cost is independent of the number of satellites, effectively mitigating the latency issue. Since each satellite sends and receives data simultaneously in each round, Ring AllReduce can make better use of parallelism and reduce idle time.
@@
$T_{\text{bc}}=(N-1)\cdot c_{\text{em}}.$ (22)
It is evident that the time consumed by the Ring Allreduce-based transmission scheme is independent of the number of satellites. Therefore, the algorithm scales well with the number of satellites, making it highly suitable for large-scale constellations.
V-B Satellite-Ground Communication for Embedding Vector Transmission
@@
V-C Inter-Orbit Communications for Global Head Aggregation
@@
To propose an aggregation algorithm, it is imperative first to elucidate the network topology. We model the LEO satellite constellation as a network where vertices represent satellites and edges denote ISLs. Given the stability of the intra-orbit ISLs, satellites within the same orbit can be considered as a tightly connected cluster and hold the same data because intra-orbit aggregation is done before inter-orbit aggregation. We enumerate the clusters from $1$ to $P$. There are a few links between cluster $i$ and cluster $i+1$, $\forall i\in\left\{1...P-1\right\}$. Cross-seam ISLs are not considered in the system, so there are no links between cluster $1$ and cluster $P$. Our goal is to select a path covering all orbits to achieve fast and stable transmission from cluster $1$ to cluster $P$. The definition of path $P_{s\rightarrow t}$ is the trajectory along which data is transmitted from $s$ to $t$ through certain relay satellites. For example, $\mathbbm{P}_{S_{\text{ab}}\rightarrow S_{\text{cd}}}$ signifies the path from satellite $S_{\text{ab}}$ to satellite $S_{\text{cd}}$ and can be expressed as $\left\{S_{\text{ab}},S_{x_{1}y_{1}},S_{x_{2}y_{2}},...,S_{\text{cd}}\right\}$, where $S_{x_{i}y_{i}}$ is i-th relay satellite, $S_{\text{ab}}$ is data source and $S_{\text{cd}}$ is the destination. $\mathbbm{P}_{O_{\text{a}}\rightarrow O_{\text{c}}}$ denotes the path from any satellite in orbit $a$ to any satellite in orbit $c$. To complete global aggregation, the selected path $\mathbbm{P}$ must satisfy:
@@
The total communication delay for global head aggregation is the time it takes for parameters to transmit from orbit $1$ to orbit $P$ on selected path $\mathbbm{P}_{O_{1}\rightarrow O_{P}}$, expressed as the sum of the transmission delay on each ISL belonging to the path:
@@
The solution needs to satisfy the constraint in ( 23). To find a path with minimum delay, the distance between satellites $S_{a,b}$ and $S_{c,d}$ is set to $\frac{||S_{a,b}S_{c,d}||}{r_{L_{(S_{a,b},S_{c,d})}}}$, which is positively correlated with the delay $T(S_{ab},S_{cd})$. The minimum delay problem is modeled as a shortest path problem. We run the Floyd-Warshall algorithm on the network to find the shortest path between each pair of satellites. Subsequently, the shortest path between orbit $1$ and orbit $P$ is selected, ensuring that transmitting data on this path incurs the shortest time.

### v3

@@ Abstract
Advancements in artificial intelligence (AI) and low-earth orbit (LEO) satellites have promoted the application of large remote sensing foundation models for various downstream tasks. However, direct downloading of these models for fine-tuning on the ground is impeded by privacy concerns and limited bandwidth. Satellite federated learning (FL) offers a solution by enabling model fine-tuning directly on-board satellites and aggregating model updates without data downloading. Nevertheless, for large foundation models, the computational capacity of satellites is insufficient to support effective on-board fine-tuning in traditional Satellite FL frameworks. To address these challenges, we propose a satellite-ground collaborative federated fine-tuning framework. The key of the framework lies in how to reasonably decompose and allocate model components to alleviate insufficient on-board computation capabilities. During fine-tuning, satellites exchange intermediate results with ground stations or other satellites for forward propagation and back propagation, which brings communication challenges due to the special communication topology of space transmission networks, such as intermittent satellite-ground communication, short duration of satellite-ground communication windows, and unstable inter-orbit inter-satellite links (ISLs). To reduce transmission delays, we further introduce tailored communication strategies that integrate both communication and computing resources. Specifically, we propose a parallel intra-orbit communication strategy, a topology-aware satellite-ground communication strategy, and a latency-minimalization inter-orbit communication strategy to reduce space communication costs. Simulation results demonstrate significant reductions in training time to 33% of on-board training time.
@@ I Introduction
A promising solution to this challenge is satellite federated learning (SFL), which enables satellites to collaboratively train AI models without requiring raw data to be transmitted to ground stations [10, 11, 12]. While SFL has been explored in prior studies, most existing work focuses on small-scale models. Despite the rapid development of Space-CPN, onboard computation capabilities remain significantly constrained compared to terrestrial AI clusters, making the fine-tuning of entire large models a major challenge. Satellite federated learning, as a solution for distributed training, supports collaborative training among satellites [10, 11]. To solve this problem, it is a feasible solution to decompose and deploy the computing tasks of fine-tuning large models in Space-CPN and use the distributed computation capabilities for federated fine-tuning [13, 14]. According to the distribution of computation capabilities in Space-CPN, the modules are deployed as follows: the embedding layer is deployed on satellites because satellites are data sources, the backbone network is deployed on the terrestrial server, and the classification layer generating outputs is deployed on satellites, as illustrated in Fig. 1.
In this deployment framework, the primary challenge lies in the communication process. During forward propagation, satellites compute local embedding vectors and perform intra-orbit aggregation. Notably, while raw satellite data can reach gigabyte levels in total size, the transmitted embedding vectors for each sample are only a few megabytes, significantly alleviating communication overhead. Once computed, these embedding vectors are transmitted to the ground for feature vector computation. In this process, satellite-ground communication faces some special challenges compared with transmission on the ground [14]. Firstly, due to the movement of satellites, especially LEO satellites that circle the Earth at high speeds, SGLs are dynamic and unstable. Secondly, it takes much time to access the GS for each satellite. Besides, the insufficient bandwidth of SGLs leads to slower data rates and higher latency. Then, feature vectors are sent back to satellites for output generation, which faces similar questions as embedding vector downloading. Finally, satellites update local models and aggregate a global model through ISLs. The dynamic ISLs make link selection and allocation extremely complex in terms of network management [15].
@@
TABLE I: Key parameters
data size calculcation
image size of 4800 $\times$ 2800 pixels 1.4Gb width $\cdot$ height $\cdot$ channels $\cdot$ 8 bits
embedding vector size of spectralGPT 20MB hidden sequence length $\cdot$ hidden dimension $\cdot$ 32 bits
feature vector size of spectralGPT 32kb feature dimension $\cdot$ 32 bits
Flops of embedding layer $1.13\times 10^{8}$flops patch number $\cdot$ patch size $\cdot$ embedding dimension
[…]
average communication window 10 minutes from satellite tool kit (STK)
average revisit intervals 3 hours from STK
@@ II-A Centralized Satellite FL
In centralized Satellite FL, a coordinator, such as a parameter server (PS), collects model updates from satellites. Satellites perform local training on their datasets and then send the model updates to the central coordinator. Some algorithms employ only SGLs during training, where models are trained on-board and transmitted via SGLs for global model aggregation at the PS. These algorithms are usually designed to address the challenges of time-varying and unstable SGLs. For instance, the algorithm proposed in [16] adjusted aggregation intervals based on the density of satellite-ground connections to increase the frequency of model updates. [17] proposed an over-the-air computation based scheme and optimized both uplink and downlink communications through beamforming. Considering the variations in connection time and connection density among satellites, asynchronous FL is introduced to mitigate the effects of stragglers, allowing satellites to send model updates to the server asynchronously [18, 19]. A compensation mechanism proposed in [20] addressed gradient staleness by adjusting the updates according to discrepancies.
@@ II-B Decentralized Satellite FL
II-C Existing Communication Strategy
Prior work on intra-orbit aggregation relies on simple sequential data collection, where a designated satellite aggregates intra-orbit updates before forwarding them. This leads to linear latency growth with the number of satellites, wasting bandwidth due to idle links. Existing strategies for satellite-ground links (SGLs) either ignore inter-satellite cooperation and each satellite transmits local data independently. For example, the method in  [16] adjusts aggregation intervals based on connection density but does not optimize transmission tasks across multiple SGLs, leading to underutilization of short communication windows. Decentralized SFL methods use inter-orbit ISLs but focus on packet loss compensation rather than latency minimization, such as  [27]. Gossip-based inter-orbit transmission requires more rounds to reach global consensus, increasing training epochs.
@@ III-B1 Ground Communication Network
During federated fine-tuning, training information is exchanged between satellites and GSs for global aggregation. As depicted in Fig. 2, there are $G$ GSs and a PS situated on the ground, and all GSs are connected with the PS. The PS serves as the central coordinator in this system, while the GSs act as relays that facilitate communication between the satellites and the PS. The communication links between the GSs and the PS are fixed and characterized by high-speed fabric connectivity, ensuring stable and fast transmission [31]. The bandwidth is much larger than the bandwidth of the SGLs, so data transmission between GSs and PS will not become a transmission bottleneck. To represent the characteristics of the ground communication network, we model it as a stable star network, as illustrated in Fig. 3(b), where the PS is the central node and transmission between the central node and other nodes is fast.
@@ III-B2 Satellites Communication Network
Figure 3: Satellite-ground communication network topology
@@ III-B3 Satellite-Ground Communication Network
When analyzing SGLs, we sample some time points discretely $\mathcal{T}=\left\{T_{0},T_{1},T_{2},...\right\}$ over a period. A satellite and a GS can communicate between $T_{i}$ and $T_{i+1}$ if at least one link is feasible at $T_{i}$. The interval between two adjacent time points is 1 minute. At different time points, the topology of SGLs varies. Thus, we model satellite-ground communication links as time-varying topology, as depicted in Fig. 3(c). For the coordination of multiple ground stations across orbits, a parallel transmission strategy is adopted. Each orbit can independently transmit data through the GSs within its coverage area.
@@ III-C Satellite Federated Fine-Tuning for Remote Sensing Foundation Models
[…]
@@ IV-A1 Intra-Orbit Embedding Vector Broadcast
First, each satellite $S_{p,n}$ computes embedding vectors $\bm{E}_{p,n}$ utilizing data stored locally. Then, all computed embedding vectors are broadcast within each orbit so that any satellite connected to a GS can transmit intra-orbit data to the ground for forward propagation. We visualize the connection windows between satellites and GSs throughout the day in Fig. 5 for comparison. Each connection window lasts for approximately 5105$\sim$ 105 ∼ 10 minutes. If data transmission starts at 23:30, some satellites establish their first satellite-ground communication around three hours later. So the conventional method, where each satellite only sends its local embedding vectors without coordination, wastes the communication resources of other satellites. In contrast, our proposed method allows other intra-orbit satellites to transmit data immediately without waiting.
Due to the instability of inter-orbit ISLs and the volume of intra-orbit embedding vectors, embedding vectors are not exchanged between different orbits. To optimize this process, we have designed a parallel intra-orbit broadcast strategy, detailed in Section V-B, to make full use of each satellite’s communication capabilities. After the broadcast, each satellite stores the concatenated intra-orbit embedding vectors, represented as:
@@ IV-A2 Embedding Vectors Transmission
Intra-orbit embedding vectors are transmitted from satellites to the PS in each orbit. A key challenge is that satellite-ground communication topology is time-varying, as illustrated in section III-B. To address this, we design a topology-aware satellite-ground transmission strategy in section V-C, which allows multiple orbits to independently utilize this strategy for parallel transmission. The algorithm takes into account the heterogeneity of different SGLs and allocates different transmission tasks to SGLs. Thus, the total transmission time $T_{\text{em}}$ is determined by the maximum transmission time across all orbits:
@@ IV-B3 Inter-orbit Head Aggregation
Head parameters are transmitted from orbit 1 to orbit $P$ through inter-orbit ISLs for global aggregation. To ensure efficient convergence and communication, the inter-orbit ISLs are carefully selected for transmission between two adjacent orbits. A path selection algorithm is proposed in section V-D, which can adapt to dynamic inter-orbit ISLs and automatically select the optimal path to minimize transmission latency.
@@ IV-B4 Global Head Broadcast
Finally, the global model is transmitted back from orbit $P$ to orbit 111 via inter-orbit ISLs. Within each orbit, the satellite receiving the global head then disseminates it to other satellites for the next round of training, where satellites continue to train using their local data. This step ensures that every satellite has access to the latest version of the global model. Notably, this can be efficiently achieved by simply reversing the direction of transmission during inter-orbit head aggregation.
@@
[…]
for synchronous round r=0,…,R  do
@@ V-A Unified Optimization Framework
The proposed communication strategies (Ring AllReduce, topology-aware scheduling, latency-minimal inter-orbit routing) can be unified under a joint communication-computation optimization framework with the objective of minimizing total training latency, defined as:
$\min_{\mathbb{P},\mathbf{x},\mathbf{y}}\quad$ $T_{\text{total}}=T_{\text{comp}}+T_{\text{comm}}$
s.t. $\sum_{n=1}^{N}x_{n}=1$
$x_{n}=\frac{1}{N}$
$\sum\mathbf{y}_{l}\leq C_{l}$
$\mathcal{O}(\mathbb{P})=\{0,1,...,P\}$
$T_{\mathbb{P}}=\sum(T_{p}+T_{q})$
[…]
V-B Intra-Orbit Communication for Partial Head Aggregation
@@
Figure 6: Ring Allreduce based intra-orbit communication. Take N=3 as an example. Satellites $\left\{S_{p,1},S_{p,2},S_{p,3}\right\}$ travel in orbit $p$. In the scatter-reduce phase, data on satellite $S_{p,i}$ is split into 333 chunks, i.e., $C_{1}^{i},C_{2}^{i}$ and $C_{3}^{i}$. At communication round 111, $S_{p,1},S_{p,2},S_{p,3}$ transmit $C_{1}^{1},C_{2}^{2},C_{3}^{3}$ respectively. So each satellite only transmits $\frac{1}{3}$ data 222 times in the scatter-reduce phase and gather phase. The total time is $2\times 2\times\frac{1}{3}\times T$, where $T$ is the time to transmit all data on a satellite.
V-B1 Scatter-Reduce Phase
@@ V-B2 Gather Phase
In the first and second phases of Ring Allreduce, each satellite receives data $N-1$ times, with each transmission involving $\frac{\operatorname{size}(data)}{N}$ bits of data, where $\operatorname{size}(data)$ denotes the total number of bits for size. Thus, when data rate is $r$, the total time is given by:
$T_{\text{intra}}=2\cdot(N-1)\cdot\frac{\operatorname{size}(\text{head})}{N\cdot r}<2\cdot\operatorname{size}(\text{head}).$ (21)
Ring AllReduce has a constant communication complexity, meaning the communication cost is independent of the number of satellites, effectively mitigating the latency issue. Since each satellite sends and receives data simultaneously in each round, Ring AllReduce can make better use of parallelism and reduce idle time. It is evident that the time consumed by the Ring Allreduce-based transmission scheme is independent of the number of satellites. Therefore, the algorithm scales well with the number of satellites, making it highly suitable for large-scale constellations.
@@
$T_{\text{bc}}=\frac{(N-1)\cdot c_{\text{em}}}{r}$ (22)
, where $c_{em}$ is the size of embedding vectors.
V-C Satellite-Ground Communication for Embedding Vector Transmission
@@
V-D Inter-Orbit Communications for Global Head Aggregation
@@
To propose an aggregation algorithm, it is imperative first to elucidate the network topology. We model the LEO satellite constellation as a network where vertices represent satellites and edges denote ISLs. Given the stability of the intra-orbit ISLs, satellites within the same orbit can be considered as a tightly connected cluster and hold the same data because intra-orbit aggregation is done before inter-orbit aggregation. We enumerate the clusters from 111 to $P$. There are a few links between cluster $i$ and cluster $i+1$, $\forall i\in\left\{1...P-1\right\}$. Cross-seam ISLs are not considered in the system, so there are no links between cluster 111 and cluster $P$. Our goal is to select a path covering all orbits to achieve fast and stable transmission from cluster 111 to cluster $P$. The definition of path $P_{s\rightarrow t}$ is the trajectory along which data is transmitted from $s$ to $t$ through certain relay satellites. For example, $\mathbbm{P}_{S_{\text{ab}}\rightarrow S_{\text{cd}}}$ signifies the path from satellite $S_{\text{ab}}$ to satellite $S_{\text{cd}}$ and can be expressed as $\left\{S_{\text{ab}},S_{x_{1}y_{1}},S_{x_{2}y_{2}},...,S_{\text{cd}}\right\}$, where $S_{x_{i}y_{i}}$ is i-th relay satellite, $S_{\text{ab}}$ is data source and $S_{\text{cd}}$ is the destination. $\mathbbm{P}_{O_{\text{a}}\rightarrow O_{\text{c}}}$ denotes the path from any satellite in orbit $a$ to any satellite in orbit $c$. To complete global aggregation, the selected path $\mathbbm{P}$ must satisfy:
@@
The total communication delay for global head aggregation is the time it takes for parameters to transmit from orbit 111 to orbit $P$ on selected path $\mathbbm{P}_{O_{1}\rightarrow O_{P}}$, expressed as the sum of the transmission delay on each ISL belonging to the path:
@@
The solution needs to satisfy the constraint in (23). To find a path with minimum delay, the distance between satellites $S_{a,b}$ and $S_{c,d}$ is set to $\frac{||S_{a,b}S_{c,d}||}{r_{L_{(S_{a,b},S_{c,d})}}}$, which is positively correlated with the delay $T(S_{ab},S_{cd})$. The minimum delay problem is modeled as a shortest path problem. We run the Floyd-Warshall algorithm on the network to find the shortest path between each pair of satellites. Subsequently, the shortest path between orbit 111 and orbit $P$ is selected, ensuring that transmitting data on this path incurs the shortest time.
V-E Complexity Analysis
Two algorithms are mainly used in this paper: the maximum flow algorithm and the shortest path algorithm. Specifically, the time complexity of the Edmonds-Karp algorithm for maximum flow is $O(V\cdot E^{2})$, where E denotes the number of edges and V represents the number of nodes. In the context of satellite-ground transmission, V is the sum of the number of ground stations and the number of satellites connected to these ground stations, and E is the number of satellite-ground links. For the Floyd-Warshall algorithm for shortest paths, its complexity is $O(V^{3})$. Routing tasks, such as determining the shortest path for inter-orbit communication (e.g., using the Floyd-Warshall algorithm to minimize latency), can be performed on the ground server, similar to routing algorithms in  [39, 40], which have significantly stronger computational power compared to on-board satellite systems. These routing results are then transmitted back to the satellites along with feature vectors or embedding vectors. Centralizing complex routing on the ground with robust computing capabilities makes the overall system feasible, given the complexity of $O(V\cdot E^{2})$ and $O(V^{3})$.